OpenStudy (gabylovesyou):

@Reaper534 Box A has a volume of 48 cubic meters. Box B is similar to Box A. To create Box B, Box A's dimensions were doubled. What is the volume of Box B? 24 m3 96 m3 < ------- ??? is it this one 288 m3 384 m3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nonono

OpenStudy (anonymous):

48x4 = ?

OpenStudy (gabylovesyou):

192

OpenStudy (anonymous):

Oh wait i read that wrong lol

OpenStudy (gabylovesyou):

lol cx

OpenStudy (anonymous):

Hello sory i was doing something ok lets see

OpenStudy (gabylovesyou):

lol dont worry

OpenStudy (anonymous):

Grrr sorry this lagg

OpenStudy (anonymous):

Okay, well we know to calculate volume of box A that the formula is: Va=l×w×h

OpenStudy (anonymous):

sorry thats Va=(l)(w)(h)

OpenStudy (anonymous):

the question says: "To create box B, box A's dimensions were doubled." Vb=l(2) x w(2) xh(2)

OpenStudy (anonymous):

or Vb= 8( l x w x h)

OpenStudy (gabylovesyou):

so its D ?

OpenStudy (anonymous):

yes

OpenStudy (gabylovesyou):

my god this lag. thanks! (。♥‿♥。)

OpenStudy (anonymous):

np haha